# Leveraging Skills from Unlabeled Prior Data for Efficient Online Exploration

### v2

@@ H.3 D4RL Play Dataset
Figure 14: Performance of our method on the play datasets. Ours outperforms all baselines, similar to the results on the diverse datasets (Figure 10). We average over 4 seeds.
Appendix I OGBench Results
@@
In this section, we include the full results on individual tasks for each of the four OGBench domains (HumanoidMaze, Cube, Scene, and AntSoccer) (Park et al., 2024a).
I.1 HumanoidMaze
@@
As shown in Figure 15, our method substantially outperforms all prior methods on the difficult HumanoidMaze environment. It is the only method to achieve nonzero return on the more difficult large and giant mazes, and performs approximately four times better than the next best baseline, Online w/ Traj. Skills, on the medium environments. These results show that on difficult, long horizon tasks, using offline data during online learning is essential for strong exploration performance.
@@
Figure 15: Normalized return on six HumanoidMaze tasks. Our method is the only method that solves the task with more than 50% success rate. All baselines either completely fail or only achieve less than 20% success rate on easier mazes (Online w/ Traj. Skill on medium-navigate and medium-stitch). We average over 8 seeds.
I.2 Cube and Scene
@@
As shown in Figure 16, Ours matches or outperforms the next best baseline on 11 of the 15 tasks. The novel baseline that we introduce HILP w/ Offline Data which also uses offline data for skill pretraining and online learning outperforms Ours on four of the Scene tasks, which further shows how using offline data twice is critical. Additionally, on two of the difficult Double Cube manipulation tasks, Ours is the only method to achieve nonzero reward. Non-skill based methods ExPLORe and Diffusion BC + JSRL performs reasonably well on the easier Cube Single and Scene task 1, but struggle to achieve significant return on the more difficult tasks, which shows how extracting structure from skills is critical for solving challenging tasks.
@@
Figure 16: Normalized return on individual tasks in Cube and Scene domains. Cube has 10 tasks in total (5 on Cube-Single and 5 on Cube-Double). Scene has 5 tasks in total. We average over 8 seeds.
I.3 AntSoccer
@@
As shown in Figure 17, Ours matches or outperforms all baselines on AntSoccer Medium, and achieves higher final performance than all baselines on AntSoccer Arena. We also see that on both tasks, Ours and HILP w/ Offline Data outperform Online w/ Traj. Skills and Online w/ HILP Skills, which demonstrates the importance of using offline data twice to accelerate online exploration.
@@
Figure 17: Normalized return on individual tasks in the AntSoccer domain. We average over 4 seeds.
Appendix J Sensitivity to Offline Data Qualities
@@
To provide more insights on how different offline data qualities affect the performance of our method, we perform additional analysis on the AntMaze-Large environment.
@@
Figure 18: Performance comparison on AntMaze-Large with different offline datasets (top-right goal). Expert: collected by a non-noisy expert policy; Navigate: collected by a noisy expert policy that randomly navigates the maze from OGBench (Park et al., 2024a); Stitch: collected by the same noisy expert policy but with much shorter trajectory length (also from OGBench (Park et al., 2024a)); Explore: random exploratory trajectories collected by moving the ant in random directions re-sampled every 10 environment steps, with large action noise added (also from OGBench (Park et al., 2024a)).
J.1 Expert Data to Random Exploratory Data
@@
As expected, the baseline ExPLORe shows a gradual performance degradation from Expert to Navigate, to Stitch, and to Explore. All skill-based methods (including our method) fail completely on Explore. This is to be expected because the Explore dataset contains too much noise and the skills extracted from the dataset are likely very poor and meaningless. The high-level policy then would have trouble composing these bad skills to perform well in the environment. On Navigate and Stitch, our method outperforms other baselines, especially on the more challenging Stitch dataset where it is essential to stitch shorter trajectory segments together. On Expert, all methods perform similarly with ExPLORe doing slightly better. We hypothesize that this is because with the expert data, online learning does not require as much exploration, and skill-based methods are mostly beneficial when there is a need for structured exploratory behaviors. Since the expert dataset has a very different distribution (much narrower) than the others, we performed a sweep over the KL coefficient in VAE training over {0.01, 0.05, 0.1, 0.2, 0.4, 0.8}, and found that 0.2 performed best, so we used these skills for both Ours and Online w/ Traj. Skills on this dataset.
J.2 Data Corruptions
@@
We report the performance on both settings in Figure 19. For the Insufficient Coverage setting, our method learns somewhat slower than the full data setting, but can still reach the same asymptotic performance, and outperforms or matches all baselines in the same data regime throughout the training process. For the 5% Data setting, our method also reaches the same asymptotic performance as in the full data regime, and outperforms or matches all baselines throughout training. The gap between Ours and baseline performance (in particular, ExPLORe) is smaller than in the full data regime, which is to be expected as we have less data to learn the prior skills, so the skills are likely not as good. Overall, among the top performing methods in the AntMaze domain, our method is the most robust, consistently outperforming the other baselines that either do not use pre-trained skills (ExPLORe) or do not use the offline data during online learning (Online w/ Trajectory Skills) in these data corruption settings.
Figure 19: Data corruption ablation on state-based antmaze-large. Top: The success rate of different methods on these data corruption settings. Bottom: Visualization of the data distribution for each corruption setting. We experiment with two data corruption settings. Our method performs worse than the full data setting but still consistently outperforms all baselines.

### v4

@@ I.3 D4RL Play Dataset
Figure 15: IQM normalized return of our method on the play datasets. SUPE outperforms all baselines, similar to the results on the diverse datasets (Figure 11). We average over 4 seeds.
Appendix J OGBench Results
@@
In this section, we include the full results on individual tasks for each of the five OGBench domains (humanoidmaze, cube-single, cube-double, scene, and antsoccer(Park et al., 2024a).
J.1 HumanoidMaze
@@
As shown in Figure 16, our method substantially outperforms all prior methods on the difficult humanoidmaze environment. It is the only method to achieve nonzero return on the more difficult large and giant mazes, and performs approximately four times better than the next best baseline, Trajectory Skills, on the medium environments. These results show that on difficult, long horizon tasks, using offline data during online learning is essential for strong exploration performance.
@@
Figure 16: IQM normalized return on six humanoidmaze tasks. SUPE is the only method that obtain more than 0.5 IQM normalized return. All baselines either completely fail or only achieve less than 0.2 IQM normalized return on easier mazes (e.g., Trajectory Skills on medium-navigate and medium-stitch). We average over 8 seeds.
J.2 Cube and Scene
@@
As shown in Figure 17, SUPE matches or outperforms the next best baseline on 12 of the 15 tasks. The novel baseline that we introduce SUPE (HILP) which also uses offline data for skill pretraining and online learning outperforms SUPE on three of the scene tasks, which further shows how using offline data twice is critical. Additionally, on one of the difficult cube-double manipulation tasks, SUPE is the only method to achieve nonzero reward. Non-skill based methods ExPLORe and DBC+JSRL performs reasonably well on the easier cube-single and scene-task1, but struggle to achieve significant return on the more difficult tasks, which shows how extracting structure from skills is critical for solving challenging tasks.
@@
Figure 17: IQM normalized return on individual tasks in cube and scene domains. cube has 10 tasks in total (5 on cube-single and 5 on cube-double). scene has 5 tasks in total. We average over 8 seeds.
J.3 AntSoccer
@@
As shown in Figure 18, SUPE (HILP) outperforms all baselines on both antsoccer-arena and antsoccer-medium. We also see that on both tasks, SUPE and SUPE (HILP) outperform Trajectory Skills and HILP Skills, which demonstrates the importance of using offline data twice to accelerate online exploration.
@@
Figure 18: IQM normalized return on individual tasks in the antsoccer domain. We average over 8 seeds.
Appendix K Sensitivity to Offline Data Quality
@@
To provide more insights on how the quality of the offline data affects the performance of our method, we perform additional analysis on the antmaze-large environment.
@@
Figure 19: Performance comparison on antmaze-large-top-right with different offline datasets. Expert: collected by a non-noisy expert policy; Navigate: collected by a noisy expert policy that randomly navigates the maze from OGBench (Park et al., 2024a); Stitch: collected by the same noisy expert policy but with much shorter trajectory length (also from OGBench (Park et al., 2024a)); Explore: random exploratory trajectories collected by moving the ant in random directions re-sampled every 10 environment steps, with large action noise added (also from OGBench (Park et al., 2024a)). We average over 4 seeds.
K.1 Expert Data to Random Exploratory Data
@@
As expected, the baseline ExPLORe shows a gradual performance degradation from Expert to Navigate, to Stitch, and to Explore. All skill-based methods (including our method) fail completely on Explore. This is to be expected because the Explore dataset contains too much noise and the skills extracted from the dataset are likely very poor and meaningless. The high-level policy then would have trouble composing these bad skills to perform well in the environment. On Navigate and Stitch, our method matches or outperforms other baselines, especially on the more challenging Stitch dataset where it is essential to stitch shorter trajectory segments together. On Expert, ExPLORe performs slightly better than all other methods. We hypothesize that this is because with the expert data, online learning does not require as much exploration, and skill-based methods are mostly beneficial when there is a need for structured exploratory behaviors. Since the expert dataset has a very different distribution (much narrower) than the others, we performed a sweep over the KL coefficient in VAE training over {0.01, 0.05, 0.1, 0.2, 0.4, 0.8}, and found that 0.2 performed best, so we used these skills for both SUPE and Trajectory Skills on this dataset.
K.2 Data Corruptions
@@
We report the performance on both settings in Figure 20. For the Insufficient Coverage setting, our method learns somewhat slower than the full data setting, but can still reach the same asymptotic performance, and outperforms or matches all baselines in the same data regime throughout the training process. For the 5% Data setting, our method also reaches the same asymptotic performance as in the full data regime, and outperforms or matches all baselines throughout training. Overall, among the top performing methods in the antmaze domain, our method is the most robust, consistently outperforming the other baselines that either do not use pre-trained skills (ExPLORe) or do not use the offline data during online learning (Trajectory Skills) in these data corruption settings.
Figure 20: Data corruption ablation on state-based antmaze-large. Top: The success rate of different methods on these data corruption settings. Bottom: Visualization of the data distribution for each corruption setting. We experiment with two data corruption settings. Our method performs worse than the full data setting but still consistently outperforms all baselines. We average over 8 seeds.